OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

OpenStudy (helder_edwin):

what is the slope of the line through the points u r given?

OpenStudy (nali):

use the formula y1-y2 = m(x1-x2)

OpenStudy (nali):

and find m

OpenStudy (lgbasallote):

\[y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] \[y - 2 = \frac{6-2}{-9 - 3} (x-3)\] \[y - 2= \frac{4}{-12} (x-3)\] \[y - 2 = -\frac 13 (x-3)\] \[3y - 6 = x - 3\] simplify this now

OpenStudy (nali):

 m=-3